Find the value of x-N^x=x , where N belongs to natural number greater than 1.

Grade:11

1 Answers

Saurabh Singh
6 years ago
I think your question is wrong as x-N^x=xCan be written asx-x=N^x=>N^x=0Hence there is no value of x for which N is a natural number greater than one .So please check your question and ask again !